OpenStudy (girlgoyleh.):

Someone please help!! The system of equations is coincident. What are the missing values? -2x+5y=-12 -6x+? y=?

OpenStudy (girlgoyleh.):

@haleyelizabeth2017

Directrix (directrix):

If the system is coincident, then the two lines have the same equation.

OpenStudy (haleyelizabeth2017):

So what do you multiply the top one by to make it where -2x equals -6x?

Directrix (directrix):

-2x+15y=-12 -6x+__ y=__ ========== Multiply each member of the first equation by 3.

OpenStudy (girlgoyleh.):

3

OpenStudy (girlgoyleh.):

-2x3=-6 so i do 15x3?

OpenStudy (haleyelizabeth2017):

So what is 15*3?

OpenStudy (girlgoyleh.):

oh wait! its actually a 5, not 15 sorry

OpenStudy (haleyelizabeth2017):

-12*3? That's all you have to do :)

OpenStudy (haleyelizabeth2017):

Yep! :D so what did you get?

OpenStudy (girlgoyleh.):

-2x+5y=-12 -6x+? y=? that is the right equation. so its 15 for the first one, and -36 for the second one

OpenStudy (haleyelizabeth2017):

You can just leave the 15 because that's what you get lol

OpenStudy (haleyelizabeth2017):

Correct! :D

OpenStudy (girlgoyleh.):

thanks!!!

OpenStudy (haleyelizabeth2017):

no problem!